OpenStudy (anonymous):

how do I solve 3/2x+1/2=10/4x

OpenStudy (anonymous):

how do I do that

OpenStudy (anonymous):

\[\frac{3}{2x}+\frac{1}{2}=\frac{10}{4x}\]\[6+2x=10\]\[2x=4\]\[x=2\]

OpenStudy (anonymous):

thank @twvogels is it okay if you help me with 2 more like this

OpenStudy (anonymous):

@RadEn can you help me please

OpenStudy (raden):

what you question ?

OpenStudy (raden):

* your

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

\[x/4-x/12=1/2\]

OpenStudy (raden):

well, this is one of way to solve the equation above. x/4 - x/12 = 1/2 first, the LCM of (4,12,2) = 12 so, times by 12 on the left side and right side then simplify, get 12 (x/4 - x/12) = 12 * 1/2 3x - x = 6 2x = 6 then divided by 2 on both sides, get 2x/x = 6/2 x = 3

OpenStudy (anonymous):

thanks I have one more can you help me with it please @RadEn

OpenStudy (anonymous):

\[x+\frac{ 12 }{ x}=7\]

OpenStudy (raden):

no problem. please, post it

OpenStudy (anonymous):

I just did

OpenStudy (raden):

ok, first multiply by x on both side then simplify get x + 12/x = 7 x(x + 12/x) = 7 * x x^2 + 12 = 7x subtract by 7x on both side, get x^2 - 7x + 12 = 0 factor out to make the zeroes : (x - 3)(x - 4) = 0 then set up all factor be zero, then solve for x : x-3 =0 get x = 3 x-4 =0 x = 4 so, as the solution = {3,4}